OpenStudy (anonymous):

4x + 3y = 9 in slope-intercept form

OpenStudy (alexandervonhumboldt2):

first subtract 4x from both sides

OpenStudy (alexandervonhumboldt2):

4x-4x+3y=9-4x

OpenStudy (dayakar):

divide each term with 9

OpenStudy (alexandervonhumboldt2):

simplify 3y=9-4x

OpenStudy (alexandervonhumboldt2):

divide both sides by 3

OpenStudy (alexandervonhumboldt2):

3y/3=(9-4x)/3 y=-4/3x+(9/3)=?

OpenStudy (poisoned_ruby):

4x + 3y = 9 -4x -4x 3y= -4x + 9 /3 /3 y= -4/3x + 3

OpenStudy (anonymous):